OpenStudy (anonymous):

Find the limit of f(x) as (a) x approaches -infinity, (b) x approaches infinity, (c) x approaches 0 from the left-hand side, and (d) x approaches 0 from the right-hand side. **Function will be added momentarily**

OpenStudy (anonymous):

\[f(x)=\] \[(x-2)/(x-1), x \le0\] and \[1/x^2, x>0\]

OpenStudy (anonymous):

as \(x\to \infty\) you are looking at \(\frac{1}{x^2}\) so your job is to compute \[\lim_{x\to \infty}\frac{1}{x^2}\] do you know that one?